OpenStudy (anonymous):

Find the value of x to the nearest tenth remember SOH CAH TOA. Make sure it in degree mode

OpenStudy (anonymous):



OpenStudy (jim766):

which one should you use? sin, cos or tan?

OpenStudy (anonymous):

im taking a guess tan ?

OpenStudy (jim766):

what sides have numbers next to them opp, adj or hyp?

OpenStudy (jim766):

hyp is always across from right angle opp is accross from the angle x adj is next to angle x (but not the hyp)

OpenStudy (anonymous):

90

OpenStudy (jim766):

what side has the length of 10? opp, adj or hyp?

OpenStudy (anonymous):

10 and 8

OpenStudy (jim766):

the side 10 is the hyp the side 8 is the opp once you identify the sides, you will know what ratio to use. SOH, CAH, TAO which one use hyp and opp?

OpenStudy (anonymous):

ok who do u do that

OpenStudy (anonymous):

how

OpenStudy (jim766):

SOH - sin is opp/hyp CAH - Cos is adj/hyp TAN - Tan is opp/adj

OpenStudy (jim766):

since this problem involve the opp and the hyp you should use sin Since you are finding an angle, use sin^-1 make sure you calculator is set to degrees (mode button) type 2nd sin (8/10) press enter

OpenStudy (anonymous):

I not have a caluator

OpenStudy (jim766):

you will need one for this type of problem got a cell phone? with calc. if so hold it side ways

OpenStudy (anonymous):

ok now wat do I d o

OpenStudy (jim766):

what are you using phone or calc

OpenStudy (anonymous):

is this the answer 0.0279243606

OpenStudy (jim766):

nope

OpenStudy (jim766):

how did you get it

OpenStudy (anonymous):

ok how about this

OpenStudy (anonymous):

you would be using sin plug this in \[\sin(x)=8/10\] then do 8/10 (gives you .8) then put .8 into your calculator and press \[\sin^{-1} \] you should get 53.1 degrees

OpenStudy (anonymous):

if you're using your phone (iPhone) just plug in .8 then press 2nd sin